r Question 1.44 Hortzontal analysie of comparaive fnnancial stataments Includes O aiaustion of dolar amount chnges and peroetage dhanges from the previous to the cumont year r Question 1.45 Current year Preceding year $325,000 250,000 What le the percentage Increase In satee from trhe preceding year to trhe current year? 070% Question 1.46 Based on the Tollowing data for the current year, what is the Inventory tamover? Sales on account during year Cost of goods sold during year Accounts receivable, beginning of year Accounts receivable, end of year Inventory, beginning of year Inventory, end of year $700,000 270,000 45,000 35,000 90,000 110,000 o 3o O 27 Question 1.47 Based on the tollowing data, what le the accounts recelivable turnover? Sales on account during year Cost of goods sold during year Accounts receivable, beginning of year Accounts receivable, end of year Inventory, beginning of year Inventory, end of year $700.000 270,000 45,000 35,000 90,000 110,000 o 175 o 28
Answer #1

1.44 (c)

Horizontal analysis is comparison of ratios or items over accounting period.To make comparison one year is selected as base year.

1.45 (c)

Percentage increase = change /preceding year sale*100

=75000/250000*100

=30%

1.46 (c)

Inventory turnover ratio = Cost of goods sold/Average inventory*100

=270000/100000

=2.7

Note Average inventory = Beginning inventory +Ending inventory/2

=90000+110000/2

=100000

1.47 (a)

Accounts receivable ratio = Sales/Average receivable*100

= 700000/40000

= 17.5

Note Average receivable= 45000+35000/2=40000